40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8885     Entry time: Fri Jul 19 23:35:52 2013
Author: Koji 
Type: Summary 
Category: LSC 
Subject: TRY DC locking 

The Y arm was locked with the TRY DC signal.

The handing off process is too complicated because there is no path from ALS to the LSC error.


 The TRY DC error signal & the gain determination

- The error signal was produced by the operation 1/SQRT(TRY) - OFFSET. The initial offset was -5.

- The sign of the TRY DC error signal depends on which side of the resonance the arm is.
  By looking at the strip chart, I determined that the sign is opposite of the ALS.
  The ALS had the gain of -25, so the TRY control gain was to be positive.

- From the strip chart on the previous entry , the slope difference between the PDH error and the TRY DC error was x500.
  The arm control with POY11 PDH had the gain of 0.2. So the target gain for the TRY DC was determined to be +100.

Handing off

- The arm was stabilized by ALS. The ALS gain was -25 with FM2/3/5/6/7/10

- YARM configuration: no trigger / no FM trigger / gain =+0 / FM5 ON / OFFSET -5

- Start handing off:
  YARM: Turned up the gain to +50

- ALS: Turned off FM6/7

- YARM: Turned on FM6/7

- ALS: Turned off FM2

- YARM: Turned on FM4

- ALS: Turned off FM3/10

- YARM: Turned on FM2/3/8/9 ON

- ALS: Reduced the gain to -15

- YARM: Increased the gain to +70

- ALS: Reduced the gain to 0

- YARM: Increased the gain to +100

HANDING OFF - DONE

Changing the offset

The offset of -5 gave the TRY of <0.1.

The detuning was reduced by giving the offset of -4. TRY went up to ~.1

The offset of -3 made TRY 0.13

The offset of -2 made TRY 0.25

The offset of -1.5 made TRY 0.4. And the arm could not be held by this error signal anymore.

 

 

Attachment 1: TRY_locking.png  28 kB  | Hide | Hide all
TRY_locking.png
ELOG V3.1.3-